# Provably Convergent Federated Trilevel Learning

### 19 Dec 2023 v1

@@ 1 Introduction
2 Related Work
2.1 Trilevel Optimization
@@ 2.2 Polyhedral Approximation
3 Asynchronous Federated Trilevel Learning
@@
3.1 Hyper-Polyhedral Approximation
@@
3.2 Asynchronous Federated Algorithm
@@
3.3 Refining Hyper-polyhedral Approximation
In this section, a novel $\mu$-cut is proposed, which can be utilized for non-convex ($\mu$-weakly convex) optimization problem and thus is more general than the traditional cutting plane designed for convex optimization (Jiao et al. 2022b; Franc, Sonnenburg, and Werner 2011). We demonstrate that the proposed $\mu$-cuts are valid, i.e., the original feasible region is a subset of the polytope that forms of $\mu$-cuts in Proposition 3.3 and 3.4. Every $T_{\rm{pre}}$ iteration, the $\mu$-cuts will be updated to refine the hyper-polyhedral approximation when $t<T_{1}$, which can be divided into three steps: 1) generating new ${\rm{I}}^{\rm{st}}$ layer $\mu$-cut, 2) generating new ${\rm{II}}^{\rm{nd}}$ layer $\mu$-cut, 3) removing inactive $\mu$-cuts.
@@ 4 Discussion
$\!{\!||\boldsymbol{z}_{i}^{t}\!-\!\boldsymbol{z}_{i}^{t-k}|{|^{2}}\!\leq\!\tau{k_{1}}\vartheta},\forall i,{\sum\nolimits_{l}\!||{\lambda_{l}^{t}}\!-\!{\lambda_{l}^{t-k}}|{|^{2}}\!\leq\!\tau{k_{1}}\vartheta},{1\!\leq\!k\!\leq\!\tau},\end{array}$ (29)
where $k_{1}>0$ is a constant. Detailed discussions about Assumption 4.3 and 4.4 are provided in Appendix I.
Theorem 4.5
(Iteration Complexity) Suppose Assumption 4.3 and 4.4 hold, we set the step-sizes as ${\eta_{\boldsymbol{x}_{i}}}\!=\!{\eta_{\boldsymbol{z}_{i}}}\!=\!\frac{2}{{L+{{\eta_{\lambda}}}M{L^{2}}+{{\eta_{\boldsymbol{\theta}}}}N{L^{2}}+8(\frac{{M\gamma{L^{2}}}}{{{{\eta_{\lambda}}}{\underline{c}_{1}}^{2}}}+\frac{{N\gamma{L^{2}}}}{{{{\eta_{\boldsymbol{\theta}}}}{\underline{c}_{2}}^{2}}})}},\forall i$, ${\eta_{\boldsymbol{\theta}}}\leq\frac{2}{{L+2c_{2}^{0}}}$ and ${{\eta_{\lambda}}}\!<\!\min\{\frac{2}{{L+2c_{1}^{0}}},\frac{1}{{30\tau{k_{1}}N{L^{2}}}}\}$. For a given $\epsilon$, we have:
@@ 5 Experiment
5.1 Distributed Robust Hyperparameter Optimization
@@
Table 1: Experimental setting in distributed robust hyperparameter optimization and distributed domain adaptation for pretraining &\&& finetuning.
@@
5.2 Distributed Domain Adaptation

### 22 Jan 2024 v2

@@ Introduction
Related Work
Trilevel Optimization
@@
Asynchronous Federated Trilevel Learning
@@
Hyper-Polyhedral Approximation
@@
Asynchronous Federated Algorithm
@@
Refining Hyper-polyhedral Approximation
In this section, a novel $\mu$-cut is proposed, which can be utilized for non-convex ($\mu$-weakly convex) optimization problem and thus is more general than the traditional cutting plane designed for convex optimization (Jiao et al. 2022b; Franc, Sonnenburg, and Werner 2011). We demonstrate that the proposed $\mu$-cuts are valid, i.e., the original feasible region is a subset of the polytope that forms of $\mu$-cuts in Proposition 1 and 2. Every $T_{\rm{pre}}$ iteration, the $\mu$-cuts will be updated to refine the hyper-polyhedral approximation when $t<T_{1}$, which can be divided into three steps: 1) generating new ${\rm{I}}^{\rm{st}}$ layer $\mu$-cut, 2) generating new ${\rm{II}}^{\rm{nd}}$ layer $\mu$-cut, 3) removing inactive $\mu$-cuts.
@@ Discussion
$\!{\!||\boldsymbol{z}_{i}^{t}\!-\!\boldsymbol{z}_{i}^{t-k}|{|^{2}}\!\leq\!\tau{k_{1}}\vartheta},{\sum\nolimits_{l}\!||{\lambda_{l}^{t}}\!-\!{\lambda_{l}^{t-k}}|{|^{2}}\!\leq\!\tau{k_{1}}\vartheta},{1\!\leq\!k\!\leq\!\tau},\end{array}$ (29)
where $k_{1}>0$ is a constant. Detailed discussions about Assumption 1 and 2 are provided in Appendix I.
Theorem 1
(Iteration Complexity) Suppose Assumption 1 and 2 hold, we set the step-sizes as ${\eta_{\boldsymbol{x}_{i}}}\!=\!{\eta_{\boldsymbol{z}_{i}}}\!=\!\frac{2}{{L+{{\eta_{\lambda}}}M{L^{2}}+{{\eta_{\boldsymbol{\theta}}}}N{L^{2}}+8(\frac{{M\gamma{L^{2}}}}{{{{\eta_{\lambda}}}{\underline{c}_{1}}^{2}}}+\frac{{N\gamma{L^{2}}}}{{{{\eta_{\boldsymbol{\theta}}}}{\underline{c}_{2}}^{2}}})}},\forall i$, ${\eta_{\boldsymbol{\theta}}}\leq\frac{2}{{L+2c_{2}^{0}}}$ and ${{\eta_{\lambda}}}\!<\!\min\{\frac{2}{{L+2c_{1}^{0}}},\frac{1}{{30\tau{k_{1}}N{L^{2}}}}\}$. For a given $\epsilon$, we have:
@@
Distributed Robust Hyperparameter Optimization
@@
Table 1: Experimental setting in distributed robust hyperparameter optimization and distributed domain adaptation.
@@
Distributed Domain Adaptation